The parking lot of a college campus has 200 lamp posts. Each lamp post has two 1200-watt light bulbs. The lights are on an average of 10 hours per night year-round. Calculate the annual electric cost for the lights if the college pays 8.5 cents per KWHR.
1
Expert's answer
2019-04-24T09:59:18-0400

Power:


"P=(200)(2)(1200)=480000W=480kW"

Time:


"t=365(24)=8760 hr"

Energy:


"E=Pt=480(8760)=4204800kWhr"

The annual electric cost for the lights:


"C=(0.085)(4204800)=\$357408"